Answer the problem given below. pH Level In chemistry, the pH of a substance is defined as pH = -log log [H*] where H+ is the hydrogen ion concentration, in moles per liter. Find the pH level of each substance. SUBSTANCE a.) Pineapple juice b.) Hair conditioner c.) Mouthwash d.) Eggs e) Tomatoes HYDROGEN ION CONCENTRATION 1.6 x 10-⁴ 0.0013 6.3 x 107-⁷ 1.6 x 10-⁸ 6.3 x 10-⁵
1
Expert's answer
2021-10-28T18:46:01-0400

a) "pH=-\log({1.6*10^{-4}})=4-\log({1.6})=3.7959."

b) "pH=-\log(0.0013)=3-\log(1.3)=2.8861."

c) "pH=-\log(6.3*10^{-7})=7-\log(6.3)=6.2007."

d) "pH=-\log(1.6*10^{-8})=8-\log(1.6)=7.7959."

e) "pH=-\log(6.3*10^{-5})=5-\log(6.3)=4.2007."